DPS CALCULATION

Weston Corporation just paid a dividend of $2 a share (i.e., D0 = $2). The dividend is expected to grow 7% a year for the next 3 years and then at 4% a year thereafter. What is the expected dividend per share for each of the next 5 years? Round your answers to two decimal places.

D1 = $
D2 = $
D3 = $
D4 = $
D5 = $
Answer #1
D1 (2*1.07)=$2.14
D2 (2.14*1.07)=$2.29(Approx).
D3 (2.29*1.07)=$2.45(Approx).
D4 (2.45*1.04)=$2.55(Approx).
D5 (2.55*1.04)=$2.65(Approx).